8f+7-3f=32 any help please with this please? :)
Answer by askmemath(368)

8f+7-3f=32
Re-Arranging
8f-3f+7 = 32
5f+7 = 32
Subtracting 7 on both sides
5f = 32-7
5f = 25
Dividing by 5 on both sides
f = 5